Find the maximum temperature in a Carnot cycle operating in a room-temperature (25?C) environment if its Carnot efficiency is 0.2.
Answer #1

efficeincy e = 1-T2/T1

T2/T1 = 1-0.2

T2/T1 = 0.8

T1 = (25+273)/0.8

T1 = 99.5 deg C